please answer. i need it asap for my assignment, thanks . here is the problem :
right cylinder with circumference of base 24(pi)-> which is equal to 3.14 .
find the lateral and total area.


thank you so much for answering . GOD BLESS!

Answer by Theo(13342)

C = 2 * pi * r
24 = 2 * pi * r
r = (12/pi)
Lateral Area of the cylinder is equal to the area of the side of the cylinder.
This is equal to length * width of the side of the cylinder.
The circumference of the base is the length of the side of the cylinder.
The width of the lateral area of the cylinder is the height of the cylinder.
You did not provide this.
I will call it x.
The lateral area of the cylinder is equal to length * width which is equal to 24 * x.
Find x and multiply 24 by it and you will have the lateral area of the cylinder.
The total area of the cylinder is the area of the bases (it has 2 of them) plus the lateral area.
The area of the base of the cylinder is equal to pi * r^2.
This is equal to pi * (12/pi)^2 which is equal to pi * 144/pi^2 which is equal to (144/pi).
That's the area of one of the bases of the cylinder.
The cylinder has 2 bases, so you need to multiply this by 2 to get (288/pi) which is the area of all the bases of the cylinder.
The total area of the cylinder is equal to the area of the 2 bases plus the lateral area of the cylinder.
This is equal to 288/pi + (24*x).
Find x and you have your answer.
It needs to be part of your problem since the problem can't be solved without it.
If you were not given x (the height of the cylinder), then your answer is:
288/pi + (24*x).
Divide the area of the bases by pi and the answer is:
91.67324722 + (24*x).
Here's a reference on right cylinders that might be helpful for you to understand.
http://www.cliffsnotes.com/study_guide/Right-Circular-Cylinders.topicArticleId-18851,articleId-18834.html